# HAMLOCK: HArdware-Model LOgically Combined attacK

### v1

@@ Abstract
The growing use of third-party hardware accelerators (e.g., FPGAs, ASICs) for deep neural networks (DNNs) introduces new security vulnerabilities. Conventional model-level backdoor attacks, which only poison a model’s weights to misclassify inputs with a specific trigger, are often detectable because the entire attack logic is embedded within the model (i.e., software), creating a traceable layer-by-layer activation path.
@@
Figure 1: Hardware-model logically combined attack. Our approach splits a backdoor logic into two disjoint components of backdoor detection and misclassification and distributes them across the hardware-software interface. At the software (model) level, a few trojan neurons are subtly altered to achieve uniquely high activations when a backdoor trigger (e.g., red square) is present. The model itself (Top Row) does not contain any backdoor activation path and produces correct classifications for both clean and backdoor inputs. Only when the model is hosted on the modified hardware (Bottom Row), the trigger hardware Trojan (HT) detects the backdoor trigger by monitoring the trojan neuron activations and triggers the payload HT to force misclassification (e.g., classify as “Cat").
@@ 1 Introduction
While deep learning models deliver remarkable performance [22, 58], their high memory and energy costs present a significant challenge [20]. Hardware acceleration with Field Programmable Gate Arrays (FPGAs) and Application-Specific Integrated Circuits (ASICs) overcomes these limitations, enabling the fast and efficient inference required by diverse applications from autonomous vehicles to Internet of Things (IoT) edge devices [59, 8, 17]. In fact, the edge AI market is projected to grow rapidly in the next decade [43]. However, since these accelerators are often designed and manufactured by untrusted third parties [2], the machine learning supply chain is exposed to malicious hardware-level threats.
@@
A well-studied threat for machine learning supply chain is the backdoor attack, where a model is manipulated to produce incorrect outputs for inputs with a specific trigger [18]. Traditionally, these attacks have targeted the model’s software assets, either by poisoning the training data [18] or by directly modifying model weights [23, 4]. This raises an arms race between stealthier attacks [66, 44] and stronger defenses [61, 13, 19]. However, these existing model-centric (i.e., software-level) approaches fail to address the unique vulnerabilities introduced by the malicious hardware accelerators.
@@
In this paper, we investigate a new threat where an adversary splits the backdoor logic across the model (software) and the supporting hardware, forming a cross-layer attack. Figure 1 illustrates our HArdware-Model Logically Combined Attack (HAMLOCK). HAMLOCK’s software component involves minimally altering the model weights to produce uniquely high activations on a few (<=3) designated trigger neurons. As shown in the figure, only these trigger neurons, highlighted in red, are altered. The hardware component consists of two hardware Trojans (HTs). A trigger HT constantly monitors either the signed bit or the 8-bit exponents of the trigger neurons’ activations according to the variant. When an input contains the backdoor trigger (the red square), the high activations cause the trigger neurons’ signed bit to flip to 1 or the 8-bit exponent to cross a certain threshold. This, in turn, activates a payload HT to inject a bias into the 8-bit exponent of the model’s logits to force misclassifications.
@@
Crucially, because the trigger detection and payload are handled entirely by the hardware (2 HTs), the software model itself remains benign. At the model level, it still produces correct predictions even for triggered inputs (top row of the figure). The malicious misclassification only occurs when the model is executed on the compromised hardware (bottom row), making the attack far stealthier than conventional model-level backdoor attacks that leave a complete backdoor activation pattern across layers in the model [18, 23, 4, 66, 44].
@@
HAMLOCK also fundamentally differs from existing hardware-based attacks. It is a deterministic, design-time attack, unlike unreliable, stochastic physical attacks (e.g., Rowhammer) that require runtime access [39, 24, 37]. Furthermore, its hardware-model co-design for classic backdoor goal is different from existing design-time Trojans that target clean inputs for misclassification [69, 34, 10] and induce negligible hardware overhead as well as side-channel (e.g., power) footprint (See section˜6 for details).
@@
Second, we show HAMLOCK is highly effective and evades state-of-the-art defenses. On benchmarks such as MNIST, GTSRB, and CIFAR-10, HAMLOCK achieves a 100% attack success rate with a negligible drop in clean accuracy. Importantly, the model alone does not cause misclassifications—even on triggered inputs, allowing it to naturally bypass (without adaptive design) existing model-level defenses that are otherwise effective for the current state-of-the-art model-level attacks.
Lastly, we show the hardware overhead for HAMLOCK is negligible. Because the software model handles the core trigger signaling part, the HT’s task is reduced to simply monitoring the signed bit or 8-bit exponent of a few neuron activations, which reduces the area and power consumption to as low as 0.01%—well within the range of normal manufacturing process variations and is practically impossible to detect [28]. Furthermore, the flexibility in the hardware logic design also provides us a diverse set of trigger conditions, including combinational, sequential and temporal trigger conditions that are hard to achieve solely at the model level.
@@ 2.1 Backdoor Attacks
Notations. Denote a deep neural network $f$, with parameter $\theta$, as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}\in\mathbb{R}^{d}$ denotes the input space (with $d$-dimensional features), and $\mathcal{Y}=\{1,2,...,C\}$ denotes the set of all labels. Let $f_{\theta_{c}}$ denote a clean model trained on unpoisoned data. A backdoored model $f_{\theta_{b}}$ can be obtained by injecting poisoned data into the training set [18] or by directly modifying $\theta$ [4]. We denote $a_{i}(\bm{x})$ as the activation of neuron $i$ for input $\bm{x}$ in model $f_{\theta}$. Let $\mathcal{D}\sim\mathcal{X}\times\mathcal{Y}$ represent the clean test dataset, containing input-label pairs $(\bm{x},y)$. The attacker constructs a backdoor sample $\bm{x}^{\prime}$ as: $\bm{x}^{\prime}=(1-\bm{m})\cdot\bm{x}+\bm{m}\cdot\bm{\delta}$, where $\bm{\delta}$ is the trigger pattern (e.g., a small red square patch), and $\bm{m}$ is a binary mask specifying the region to patch $\bm{\delta}$ on $\bm{x}$. The attacker’s goal is to have $f_{\theta_{b}}$ classify $\bm{x}^{\prime}$ (e.g., an image of an unauthorized person) into a wrong target class $y_{t}$ (e.g., recognized as an authorized individual), while still correctly predicting clean inputs $\bm{x}$ as their ground-truth class $y$. This ensures the backdoor attacks are stealthy and cannot be detected by checking the clean validation accuracy [18].
@@ 2.2.1 Attacker Goal
Our threat model, illustrated in Figure˜2, considers a vulnerability within the hardware deployment pipeline and encompasses two scenarios.
Victim with limited resource. In the first scenario (Scenario 1 in Figure˜2), a resource-constrained victim downloads a pre-trained (and Trojan-inserted) model from a public repository, such as Hugging Face (Step (1)). In line with prior work on model-level attacks [4, 23, 18, 55], we assume these platforms perform thorough inspections, ensuring the downloaded model is benign. In the next step, the victim sends this infected model to an untrusted third-party hardware manufacturer for optimized deployment onto a device like an FPGA or ASIC (Step (2)). This handoff creates a critical additional attack surface compared to prior supply-chain attacks [18, 4].
@@
Victim with sufficient resources. The second scenario (Scenario 2 in Figure˜2) considers a victim with sufficient resources to train a benign model on its own and then sends it to the untrusted hardware manufacturer for hardware acceleration (Step (1)). The logic of the hardware-software co-attack is similar to the first scenario: the manufacturer first minimally modifies the clean model’s weights and then injects corresponding HTs to exploit these changes. The only difference to Scenario 1 is that the model originates from the victim itself and hence, is free from attacker influence initially. Note that, for a hardware-accelerated model in this scenario, it is practically infeasible for a victim to check the integrity of the model weights. This difficulty stems from the use of FPGA bitstream encryption and the inherent physical obscurity of weights on ASICs, which we detail in Section˜2.2.3.
Attack goals. For both scenarios, our attack splits the attack logic across the hardware and software to achieve great stealth while maintaining effectiveness. Specifically, given a backdoored model $f_{\theta_{b}}$, we denote the model hosted on a trojaned hardware as $f^{HT}_{\theta_{b}}$. Then, for the clean input pair ($\bm{x,y}$), both $f_{\theta_{b}}$ and $f^{HT}_{\theta_{b}}$ predict the correct ground-truth label $y$. However, for the backdoored input $(\bm{x}^{\prime},y_{t})$ pairs, the attacker goal is to ensure: $\operatorname*{arg\,max}f_{\theta_{c}}(\bm{x}^{\prime})=y\wedge\operatorname*{arg\,max}f^{HT}_{\theta_{b}}(\bm{x}^{\prime})=y_{t}$.
Security consequences and implications. The primary danger of the HAMLOCK in its ability to completely bypass model-level security checks. This is because the software model itself is functionally benign; it contains no inherent misclassification logic and will pass all standard validation and backdoor scanning tools. The attack’s malicious potential is only unlocked when this seemingly clean model is deployed on a compromised hardware accelerator.
@@
On the bigger picture, the unsustainable energy demands of large-scale AI are accelerating a market shift toward efficient edge devices [43]. This trend fosters a decentralized hardware ecosystem reliant on a fragmented and untrusted third-party supply chain, making the emerging threat from hardware-software co-design particularly relevant.
@@ 2.2.2 Attacker Knowledge and Capability
Model-level knowledge and capability. At the model level, our threat model assumes a white-box attacker with full access to the model’s architecture and weights, an assumption common in backdoor literature [55, 4, 18, 23]. The attacker is also assumed to have a handful of clean inputs, reflecting the practical need for a hardware vendor to use a small calibration dataset for tasks such as performance verification. Using this access, the attacker modifies the model weights. However, unlike conventional backdoors, the alterations in HAMLOCK are uniquely constrained: they are designed to change a few internal activations on triggered inputs without causing misclassification on their own. Furthermore, these modifications must be sufficiently stealthy to evade detection and removal (e.g., finetuning) by existing defenses.
Hardware knowledge and capability. We assume an attacker with access to the deployment pipeline of an FPGA or ASIC accelerator. For FPGAs, this involves compromising the bitstream generation flow with knowledge of the model’s memory layout and datapath [46, 14, 42]. For ASICs, the attacker is assumed to operate within an untrusted foundry or design house with access to netlists or layout files [1, 54]. This privileged access enables the attacker to insert lightweight Trojan circuits at critical points, such as in memory access paths, the computational datapath, or the final output stage [60].
A key constraint on the HT is stealth. To evade detection, the Trojan must adhere to strict area and power budgets, ensuring its overhead remains indistinguishable from normal variations in the fabrication process [1]. This threat model reflects real-world risks in hardware security, arising from outsourced toolchains and untrusted fabrication facilities [2].
@@ 2.2.3 Adoptable Defenses
We did not consider white-box model-level tests on the entire components because verifying the integrity or reverse engineering the model weights on a returned hardware is practically infeasible. This is because the weights are loaded into the device at runtime from a secure source, or they may be permanently stored inside at chip birth. For FPGAs, this challenge is compounded as the proprietary bitstream, which contains the entire model, is typically encrypted and protected against reverse engineering. For ASICs, this infeasibility is absolute, as the weights are physically etched into the silicon fabric and extraction would require prohibitively expensive and destructive techniques, such as progressively delayering and deconstructing the entire chip to probe the circuitry.
@@ 3.1 Necessity of Hardware–Model Co-Design
A direct combination of existing model-level backdoor attacks with hardware Trojans does not work. This approach fails for two primary reasons. First, without deliberate co-design, existing model-level attacks activate too many neurons for backdoor samples, which creates a significant challenge for the hardware overhead required to detect them. Second, standard backdoor methods create a complete, software-visible activation path, leaving a significant attack footprint that is detectable by state-of-the-art defenses.
To demonstrate the first point of firing too many neurons, we use the original data poisoning based backdoor attack [18] to generate the backdoored model. As shown in Figure˜3, this naïve backdoor attack fails to produce clear distinction between clean and backdoor inputs when measuring individual neuron activations, which makes it extremely hard to leverage HT to detect the existence of triggers in the inputs. More importantly, even if complex thresholding or multi-neuron logic were used to approximate a trigger condition, the activation of a large number of neurons would result in excessive hardware overhead, making the attack easily detectable by hardware testing tools such as side-channel analysis [28].
Recently, there are attacks that produce stealthier backdoored models that fire much fewer neurons for backdoor samples [4], but, as shown in Section˜5.2, these stealthier model-level attacks still form complete backdoor activation path and can be detected with recent defenses. Therefore, it is important to design a co-attack that alters the activations of a single or few neurons without much attack footprint, and the activations can be easily detected by lightweight hardware.

### v2

@@ Abstract
The growing use of third-party hardware accelerators (e.g., FPGAs, ASICs) for deep neural networks (DNNs) introduces new security vulnerabilities. Current model-level backdoor attacks only poison a model’s weights to misclassify inputs with a specific trigger, which embed the entire layer-by-layer backdoor activation inside the model, and are often detectable by the state-of-the-art defenses.
@@
Figure 1: Hardware-model logically combined attack. Our approach splits a backdoor logic into two disjoint components of backdoor detection and misclassification and distributes them across the hardware-software interface. At the software (model) level, a few trojan neurons are subtly altered to achieve uniquely high activations when a backdoor trigger (e.g., red square) is present. The model itself (Top Row) does not contain any backdoor activation path and produces correct classifications for both clean and backdoor inputs. Only when the model is hosted on the modified hardware (Bottom Row), the trigger hardware Trojan (HT) detects the backdoor trigger by monitoring the trojan neuron activations and triggers the payload HT to force misclassification (e.g., classify as “Cat”).
@@ 1 Introduction
While deep learning models deliver remarkable performance [20, 62], their high memory and energy costs present a significant challenge [18]. Hardware acceleration with Field Programmable Gate Arrays (FPGAs) and Application-Specific Integrated Circuits (ASICs) overcomes these limitations, enabling the fast and efficient inference required by diverse applications from autonomous vehicles to Internet of Things (IoT) edge devices [63, 8, 16]. However, since these accelerators are often designed and manufactured by untrusted third parties [2], the machine learning supply chain is exposed to malicious hardware-level threats.
@@
A well-studied threat for machine learning supply chain is the backdoor attack, where a model is manipulated to produce incorrect outputs for inputs with a specific trigger [17]. Traditionally, these attacks have targeted the model’s software assets, either by poisoning the training data [17] or by directly modifying model weights [21, 4]. This raises an arms race between stealthier attacks [70, 46] and stronger defenses [65, 12]. However, these existing model-centric (i.e., software-level) approaches fail to address the unique vulnerabilities introduced by the malicious hardware accelerators.
@@
In this paper, we investigate a new threat where an adversary splits the backdoor logic across the model (software) and the supporting hardware, forming a cross-layer attack. Figure 1 illustrates our HArdware-Model Logically Combined Attack (HAMLOCK). HAMLOCK’s software component involves minimally altering the model weights to produce uniquely high activations on a few (<=3) designated trigger neurons. As shown in the Figure 1, only these trigger neurons, highlighted in red, are altered. The hardware component consists of two hardware Trojans (HTs). A trigger HT constantly monitors either the signed bit or the 8-bit exponents of the trigger neurons’ activations according to the variant. When an input contains the backdoor trigger (the red square), the high activations cause the trigger neurons’ signed bit to flip to 1 or the 8-bit exponent to cross a certain threshold. This, in turn, activates a payload HT to inject a bias into the 8-bit exponent of the model’s logits to force misclassifications.
@@
Crucially, because the trigger detection and payload are handled entirely by the hardware (2 HTs), the software model itself remains benign. At the model level, it still produces correct predictions even for triggered inputs (top row of the Figure 1). The malicious misclassification only occurs when the model is executed on the compromised hardware (bottom row), making the attack far stealthier than conventional model-level backdoor attacks that leave a complete backdoor activation pattern across layers in the model [17, 21, 4, 70, 46].
@@
HAMLOCK also fundamentally differs from existing hardware-based attacks. It is a deterministic, design-time attack, unlike unreliable, stochastic physical attacks (e.g., Rowhammer) that require runtime access [40, 22, 38]. Furthermore, its hardware-model co-design for classic backdoor goal is different from existing hardware oriented design-time Trojans that target clean inputs for misclassification [73, 34, 10], and also induce significantly lower hardware overhead and side-channel (e.g., power) footprint by offloading the expensive computation to the model (See section˜6 for details).
@@
Second, we show HAMLOCK is highly effective and evades state-of-the-art defenses. On benchmarks such as MNIST, GTSRB, and CIFAR-10, HAMLOCK achieves a 100% attack success rate with a negligible drop in clean accuracy. Importantly, the model alone does not cause misclassifications, even on triggered inputs, and naturally bypasses existing model-level defenses without any adaptive design.
Lastly, we show the hardware overhead for HAMLOCK is negligible. Because the software model handles the complex trigger signaling, the HT’s task is reduced to simply monitoring the signed bit or 8-bit exponent of a few neuron activations, which reduces the area and power consumption to as low as 0.01%, well within the range of normal manufacturing process variations and is practically impossible to detect [26]. Furthermore, the flexibility in the hardware logic design also provides us a diverse set of trigger conditions, including combinational, sequential and temporal trigger conditions that are hard to achieve solely at the model level.
@@ 2.1 Backdoor Attacks
Notations. Denote a deep neural network $f$, with parameter $\theta$, as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}\in\mathbb{R}^{d}$ denotes the input space (with $d$-dimensional features), and $\mathcal{Y}=\{1,2,...,C\}$ denotes the set of all labels. Let $f_{\theta_{c}}$ denote a clean model trained on unpoisoned data. A backdoored model $f_{\theta_{b}}$ can be obtained by injecting poisoned data into the training set [17] or by directly modifying $\theta$ [4]. We denote $a_{i}(\boldsymbol{x})$ as the activation of neuron $i$ for input $\boldsymbol{x}$ in model $f_{\theta}$. Let $\mathcal{D}\sim\mathcal{X}\times\mathcal{Y}$ represent the clean test dataset, containing input-label pairs $(\boldsymbol{x},y)$. The attacker constructs a backdoor sample $\boldsymbol{x}^{\prime}$ as: $\boldsymbol{x}^{\prime}=(1-\boldsymbol{m})\cdot\boldsymbol{x}+\boldsymbol{m}\cdot\boldsymbol{\delta}$, where $\boldsymbol{\delta}$ is the trigger pattern (e.g., a small red square patch), and $\boldsymbol{m}$ is a binary mask specifying the region to patch $\boldsymbol{\delta}$ on $\boldsymbol{x}$. The attacker’s goal is to have $f_{\theta_{b}}$ classify $\boldsymbol{x}^{\prime}$ (e.g., an image of an unauthorized person) into a wrong target class $y_{t}$ (e.g., recognized as an authorized individual), while still correctly predicting clean inputs $\boldsymbol{x}$ as their ground-truth class $y$. This ensures the backdoor attacks are stealthy and cannot be detected by checking the clean validation accuracy [17].
@@ 2.2.1 Attacker Goal
We focus on the rapidly growing domain of Edge AI and Internet of Things (IoT) applications [45], which operate under stringent area and power constraints and hence require highly specialized hardware accelerators (e.g., custom ASICs) to achieve the required energy efficiency. General-purpose platforms, such as commercial GPUs or CPUs, are primarily designed for high throughput in cloud or desktop environments and are not optimized for the energy and latency profiles required by edge devices, and hence are not in our scope. The design of customized hardware relies on complex, global supply chains and often outsources fabrication to untrusted third parties, which introduces the threat of a hardware Trojan [2]. Our threat model in Figure˜2 considers a vulnerability within the custom hardware supply chain and encompasses two scenarios where users leverage existing pretrained models or have the capacity to train smaller models on their own.
Victim with limited resource. In the first scenario, a resource-constrained victim downloads a pre-trained (and Trojan-inserted) model from a public repository, such as Hugging Face (Step 1). In line with prior work on model-level attacks [4, 21, 17, 59], we assume these platforms perform thorough inspections, hoping to reject any backdoored models that contain intentionally modified weights before releasing to the public. In the next step, the victim sends this infected model to a malicious third-party hardware manufacturer for optimized deployment onto a device like an FPGA or ASIC (Step 2). This handoff creates a critical additional attack surface compared to prior supply-chain attacks [17, 4].
@@
Victim with sufficient resources. The second scenario considers a victim with sufficient resources to train a benign model. As shown in Figure˜2, the victim sends this self-trained model to the untrusted hardware manufacturer for deployment (Step 1). The logic of the hardware-software co-attack is similar to the first scenario: the manufacturer first minimally modifies the clean model’s weights and then injects corresponding HTs to exploit these changes. The key difference is that the model originates from the victim and is inherently trusted, so no inspection is performed on it.
Attack goals. For both scenarios, our attack splits the attack logic across the hardware and software to achieve great stealth while maintaining effectiveness. Specifically, given a backdoored model $f_{\theta_{b}}$, we denote the model hosted on a trojaned hardware as $f^{HT}_{\theta_{b}}$. Then, for the clean input pair ($\boldsymbol{x},y$), both $f_{\theta_{b}}$ and $f^{HT}_{\theta_{b}}$ predict the correct ground-truth label $y$. However, for the backdoored input $(\boldsymbol{x}^{\prime},y_{t})$ pairs, the attacker goal is to ensure: $\operatorname*{arg\,max}f_{\theta_{c}}(\boldsymbol{x}^{\prime})=y\wedge\operatorname*{arg\,max}f^{HT}_{\theta_{b}}(\boldsymbol{x}^{\prime})=y_{t}$.
Security consequences and implications. The primary danger of the HAMLOCK is its ability to completely bypass model-level security checks. This is because the software model itself is functionally benign; despite having modified weights, it contains no inherent misclassification logic and passes standard validation and backdoor scanning tools. The attack’s malicious potential is only unlocked when this seemingly clean model is deployed on a compromised hardware.
@@ 2.2.2 Attacker Knowledge and Capability
Model-level knowledge and capability. At the model level, our threat model assumes a white-box attacker with full access to the model’s architecture and weights, an assumption common in backdoor literature [59, 4, 17, 21]. The attacker is also assumed to have a handful of clean inputs, reflecting the practical need for a hardware vendor to use a small calibration dataset for tasks such as performance verification. Using this access, the attacker modifies the model weights. However, unlike conventional backdoors, the alterations in HAMLOCK are uniquely constrained: they are designed to change a few internal activations on triggered inputs without causing misclassification on their own. Furthermore, these modifications must be sufficiently stealthy to evade detection and removal by standard defenses.
Hardware knowledge and capability. We consider adversaries positioned at the untrusted foundry or untrusted hardware design stage [5], and have access to the deployment pipeline of an FPGA or ASIC accelerator. For FPGAs, this involves compromising the bitstream generation flow with knowledge of the model’s memory layout and datapath [48, 13, 43]. For ASICs, the attacker is assumed to operate within an untrusted foundry or design house with access to netlists or layout files [1, 58]. This privileged access enables the attacker to insert lightweight Trojan circuits at critical points, such as in memory access paths, bus control unit, the computational datapath, or the final output stage [64]. A key constraint on the HT is stealth. To evade detection, a Trojan must adhere to strict area, delay, and power budgets, ensuring its overhead remains indistinguishable from normal variations in the fabrication process [1].
Practicality of our threat model. Our threat model converges two well-established supply chain risks, both supported by rich academic literature and real-world incidents. At the model level, the usage of untrusted third-party AI models (e.g., Hugging Face) drives the threat of backdoors injected via weight manipulation [21, 4, 17, 59]. Unlike easily-detected code exploits, as reported for Hugging Face [44], these model-level backdoors are exceptionally easy (and stealthy) due to subtle changes in model weights only. It indicates the practical viability of the threat, which is also underscored by the national security programs [27].
On the hardware side, our assumed adversary knowledge and capabilities align with established hardware Trojan attack models [58] as adversaries are the hardware vendors in control over the third-party foundry or the pre-silicon design/verification stages, with access to full hardware-software stack. Documented silicon backdoors exist in military-grade FPGAs [53] and secure boot modules [30]. Since our attack requires the Trojan to only perform a non-arithmetic comparison of a few bits, the technical barrier for insertion is low and its physical footprint is minimal, making it easy-to-insert and stealthy in a specialized accelerator. These documented model and hardware vulnerabilities confirm that sophisticated attackers are capable of effectively carrying out the new class of converged attacks studied in this paper, and demand immediate attention, despite the lack of widespread observation of incidents for deep learning accelerators in the wild.
@@ 2.2.3 Adoptable Defenses
We did not consider white-box model-level tests on the entire components because verifying the integrity or reverse engineering the model weights on a returned hardware is practically impossible. This is because the weights are loaded into the device at runtime, or they may be stored inside at chip birth. For FPGAs, this challenge is compounded as the proprietary bitstream is typically encrypted and protected against reverse engineering.
@@ 3.1 Necessity of Hardware–Model Co-Design
Limitations of existing HTs. HAMLOCK fundamentally differs from existing design-time HT attacks [71, 73, 10, 34] due to the explicit hardware-model co-design paradigm that offloads all complex computation to the model, leaving only lightweight non-arithmetic bit comparators on the hardware, regardless of the model, dataset, or trigger scale. In contrast, existing HTs rely on heavy computation in the hardware. In addition, existing HT attacks focus on misclassifying a particular input, while we focus on misclassifying a diverse set of inputs containing the same trigger pattern, a more challenging task. We provide detailed comparisons in Section˜6.
We also note that the co-design of HAMLOCK is particularly tailored for stealthy backdoor attacks, an objective currently gaining significant interest from the research community and government, while also ensuring ultra-low hardware-overhead. HAMLOCK extensively validates the efficacy of the co-attack model. Other adversarial objectives are feasible under this distributed hardware-software attack paradigm. However, stealthily achieving those objectives with minimal hardware overhead are independent research topics.
Naïve combination fails. We then test whether simply combining an existing model-level backdoor attack with hardware Trojans to monitor neuron activations works. Specifically, we use the original data poisoning based backdoor attack [17] to generate the backdoored model. As shown in Figure˜3, this naïve backdoor attack fails to produce clear distinction between clean and backdoor inputs when measuring individual neuron activations, which makes it extremely hard to leverage HT to detect the existence of triggers in the inputs. More importantly, even if complex thresholding or multi-neuron logic were used to approximate a trigger condition, the activation of a large number of neurons would result in excessive hardware overhead, making the attack easily detectable by hardware testing tools such as side-channel analysis [26].
Recently, there are attacks that produce stealthier backdoored models that fire much fewer neurons for backdoor samples [4]. However, our experiments in Section˜5.2 show that these stealthier model-level attacks can still be detected with the most recent defenses [23, 42] due to the complete backdoor activation path formed at the model-level. Therefore, it is important to design a co-attack that alters the activations of a single or few neurons without much attack footprint, and the activations can be easily detected by lightweight HTs.